OpenStudy (anonymous):

Find the set of values of x for the following inequality (x ^{2}+4x-3)div(x^{2}+1)

OpenStudy (anonymous):

\[(x ^{2}+4x-3)\div(x^{2}+1)<x\]

OpenStudy (akshay_budhkar):

http://www.wolframalpha.com/input/?i=%28x2%2B4x%E2%88%923%29%C3%B7%28x2%2B1%29%3Cx this ill help you i believe:)

OpenStudy (anonymous):

ok... is good... thx, but i need the steps...

myininaya (myininaya):

\[\frac{x^2+4x-3}{x^2+1}-x<0\] \[\frac{x^2+4x-3}{x^2+1}-\frac{x(x^2+1)}{x^2+1}<0\] \[\frac{x^2+4x-3-x^3-x}{x^2+1}<0\] x^2+1>0 so we can multiply both sides by x^2+1 and the inequality still holds \[-x^3+x^2+3x-3<0\] \[x^3-x^2-3x+3>0\] since we multiply by -1 on both sides the inequality flipped now hopefully we can find some rational zeros to x^3-x^2-3x+3=y lets try factor by grouping \[x^2(x-1)-3(x-1)=(x-1)(x^2-3)=0=> x=1, x=\sqrt{3},x=-\sqrt{3}\] it works! now we have to test the intervals (-inf,-sqrt{3}) and (-sqrt{3},1) and (1,sqrt{3}) and (sqrt{3},inf) to see where y>0

myininaya (myininaya):

you just choose a number in each of those intervals and where you have y>0 then that interval is apart of your solution

OpenStudy (akshay_budhkar):

hats off to myininaya! hope i could give you more medals!

myininaya (myininaya):

i'm going to bed if neoh does not have any questions

myininaya (myininaya):

at first i was gonna try synthetic division but i seen that factor by grouping might work and i was happy lol

OpenStudy (anonymous):

---- +++ --- +++ -----|-----|-----|---- -sqrt{3} sqrt{3} 1 so {x:-sqrt{3}<x<sqrt{3} or x>1} is the ans??

myininaya (myininaya):

wait 1 comes first then sqrt{3}

myininaya (myininaya):

f(-2)=-3 f(0)=3 f(1.5)=-.375 f(4)=39 so (-sqrt{3},1) U (sqrt{3},inf)

OpenStudy (anonymous):

--- +++ --- +++ -----|-----|-----|---- -sqrt{3} 1 sqrt{3} so {x:-sqrt{3}<x<1 or x>sqrt{3}} is the ans??

myininaya (myininaya):

yep!

OpenStudy (anonymous):

ok... thx for the sloving the problem.. ^^

myininaya (myininaya):

np

myininaya (myininaya):

good night